OpenStudy (anonymous):

y=(lnx^2)/(ln⁡(x^2-1)), find Y'

OpenStudy (anonymous):

\[\frac{\textrm{d}}{\textrm{d}x}\left[\frac{f(x)}{g(x)}\right] =\frac{g(x)f'(x)-f(x)g'(x)}{{\textbf{[}g(x)\textbf{]}}^{\large2}}\\ y=\frac{\ln x^{\large2}}{\ln{(x^{\large2}-1)}} \\ \begin{align} y' &= \frac{\ln(x^{\large2}-1)\cdot\dfrac{1}{\ln x^{\large2}}\cdot2x- \ln x^{\large2}\cdot\dfrac{1}{\ln (x^{\large2}-1)}\cdot2x}{(\ln(x^{\large2}-1))^{\large2}} \\ &=\dfrac{2x}{\ln(x^{\large2}-1)\cdot\ln x^{\large2}} - \dfrac{2x\ln x^{\large2}}{(\ln(x^{\large2}-1))^{\large3}} \\ &=\dfrac{x}{\ln(x^{\large2}-1)\cdot\ln x} - \dfrac{4x\ln x}{(\ln(x^{\large2}-1))^{\large3}} \\ &=\dfrac{x\ln(x^{\large2}-1))^{\large2}}{\ln x\cdot(\ln(x^{\large2}-1))^{\large3}} - \dfrac{4x(\ln x)^{\large2}}{(\ln(x^{\large2}-1))^{\large3}\cdot\ln x} \\ &=\dfrac{x\ln(x^{\large2}-1))^{\large2} - 4x(\ln x)^{\large2}} {\ln x\cdot(\ln(x^{\large2}-1))^{\large3}} \end{align} \] I don't think that I got that right after all of that mess, but you get the basic idea, with the Quotient Rule.

OpenStudy (anonymous):

yeah thanks alot

OpenStudy (campbell_st):

the error occurs in the 1st line of simplifying the derivative it should be \[\frac{\frac{2xln(x^2 -1)}{x^2} - \frac{2xln(x)}{x^2 -1}}{(\ln(x^2 -1))^2}\]

OpenStudy (anonymous):

I kinda combined 2 steps into 1. What I simplified to is what you would get if you distributed your denominator of \((g(x))^{\large2}\).

OpenStudy (campbell_st):

but you log derivatives are incorrect \[y=\log(f(x))... then ...y' = \frac{f'(x)}{f(x)}\]

OpenStudy (anonymous):

Oh shiznuggets, you're right. :b